OpenStudy (anonymous):

Please help. What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm?

OpenStudy (anonymous):

well \[s=\theta r\]

OpenStudy (anonymous):

Use the equation: \[s = r \theta\] Solve for whatever you need...

OpenStudy (anonymous):

you have s and have r

OpenStudy (anonymous):

Remember that theta is in radians.

OpenStudy (anonymous):

so \[4.189=\theta*3\]

OpenStudy (anonymous):

4.189 cm ?

OpenStudy (anonymous):

Theta... radians... not centimeters, thats not an angle.

OpenStudy (anonymous):

so to find theta you divide by three and have \[\frac{ 4.189 }{ 3 }=\theta\]

OpenStudy (anonymous):

1.396 ?

OpenStudy (anonymous):

am I correct?

OpenStudy (anonymous):

you should be